# 00footnotetext: Supplementary website: https://sites.google.com/view/hetgps

### v3

@@ 1 Introduction
[…]
@@
Algorithmic Search. To study heterogeneity in MARL settings not covered by our theoretical analysis, we develop Heterogeneity Gain Parameter Search (HetGPS), a gradient-based algorithm that optimizes parameters $\theta$ of underspecified, differentiable MARL environments via backpropagation to find configurations that maximize or minimize the empirical $\Delta R$ (we assume differentiability for training efficiency, but consider non-differentiable environments in App. P). While HetGPS can in principle optimize any differentiable environment feature to influence $\Delta R$, we use it here to explore reward structures exclusively, as a means of verifying and extending our theoretical insights. Maximizing the heterogeneity gain allows us to discover reward functions where behavioral diversity is essential. Minimizing the gain leads us to settings where homogeneous policies are sufficient.
@@
Experiments. We validate our theoretical insights, and HetGPS, in simulation, by evaluating in both single-shot and long-horizon reinforcement learning environments whose reward structure instantiates the kinds of aggregation operators studied. First, in a continuous and a discrete matrix game, we test reward structures based on all nine possible combinations of $\{\min,\mathrm{mean},\max\}$, and find that the heterogeneity gains that result from the agents’ learned policies match our theoretical predictions: concave outer aggregators and convex inner aggregators benefit heterogeneous teams. Next, we test the same aggregators in embodied, partially observable environments: Multi-goal-capture, tag, and football. We find that our theory also transfers to such long-horizon MARL settings, and show that reward structures that maximize heterogeneity are meaningful and practically useful. Finally, we find that the empirical heterogeneity gain disappears as the richness of agents’ observations is increased, recovering the finding that rich observations allow agents with identical policy networks to be behaviorally heterogeneous (Bettini et al., 2023; Leibo et al., 2019).
@@
We then turn to HetGPS. Across two parameterizable families of aggregators (Softmax and Power-Sum), we show that, despite running on embodied environments, HetGPS rediscovers the reward regimes predicted by our curvature theory to maximize the heterogeneity gain, validating both HetGPS and the connection between our theoretical insights and MARL reward design (Sec. 5).
@@ 1.1 Related Works
Behavioral Diversity in MARL. Behavioral heterogeneity, where capability-identical agents learn distinct policies, can markedly improve exploration, robustness, and reward (Bettini et al., 2023). Yet heterogeneity reduces parameter sharing and thus sample-efficiency, so a core practical question is when its benefits outweigh that cost. Existing MARL methods typically adopt one of two poles: endowing each agent with its own network, or enforcing parameter sharing so all agents follow a single policy (Gupta et al., 2017a; Rashid et al., 2020; Foerster et al., 2018; Kortvelesy & Prorok, 2022; Sukhbaatar et al., 2016). A large body of work explores the efficiency–diversity trade-off (Christianos et al., 2021; Fu et al., 2022) by interpolating between these extremes: e.g., injecting agent IDs into the observation (Foerster et al., 2016; Gupta et al., 2017a), masking different subsets of shared weights (Li et al., 2024b), sharing only selected layers (Li et al., 2021), pruning a shared network into agent-specific sub-graphs (Kim & Sung, 2023), or producing per-agent parameters with a hypernetwork (Tessera et al., 2024). Further, several methods for promoting behavioral diversity in MARL have been proposed, such as: conditioning agents’ policies on a latent representation (Wang et al., 2020), decomposing and clustering action spaces (Wang et al., 2021), dynamically grouping agents to share parameters (Yang et al., 2022), applying structural constraints to the agents’ policies (Bettini et al., 2024), or by intrinsic rewards that maximize diversity (Li et al., 2021; Jaques et al., 2019; Wang et al., 2019; Jiang & Lu, 2021; Mahajan et al., 2019; Liu et al., 2023; 2024; Li et al., 2024a). While these studies demonstrate how to obtain diversity, they presume tasks where heterogeneity is advantageous. Our work addresses the orthogonal question of when diversity is beneficial, giving a principled characterization of which reward structures create that incentive in the first place.
@@ 3 Analysis
Focusing on continuous allocations, we ask what properties of aggregators guarantee $\Delta R>0$. We draw on the concept of Schur-convexity. Schur-convex functions can be understood as generalizing symmetric, convex aggregators: every convex and symmetric function is Schur-convex, but a Schur-convex function is not necessarily convex (Roberts & Varberg, 1974; Peajcariaac & Tong, 1992). Proofs for all results are available in App. G.
@@ 4 Heterogeneity Gain Parameter Search (HetGPS)
In complex scenarios where theory might be less applicable, we study heterogeneity through algorithmic search. We consider the setting of a Parametrized Dec-POMDP (PDec-POMDP, defined in App. O). A PDec-POMDP represents a Dec-POMDP (Oliehoek et al., 2016), where the observations, transitions, or reward are conditioned on parameters $\theta$. Hence, the return obtained by the agents, $G^{\theta}(\mathbf{\pi})$, can be differentiated with respect to $\theta$: $\nabla_{\theta}G^{\theta}(\mathbf{\pi})=\frac{\partial}{\partial_{\theta}}G^{\theta}(\mathbf{\pi})$. In particular, computing this gradient in a differentiable simulator allows us to back-propagate through time and optimize $\theta$ via gradient ascent [^3] Although the same approach can train policies (Xu et al., 2022; Song et al., 2024), HetGPS instead optimizes environment parameters and policies separately, using standard zeroth-order policy-gradient methods, to avoid being trapped in local minima. The overhead of implementing HetGPS in this way is modest: it increased training time by roughly 25% in our Sec. 5 experiments compared to training on a fixed environment..
@@ 5 Experiments
To empirically ground our theoretical analysis, we conduct a three-stage experimental study in cooperative MARL. We first analyze a one-step, observation-free matrix game in which each agent allocates effort $r_{ij}$ over $M$ tasks, and consider reward structures defined by aggregator pairs $U,T\in\{\min,\mathrm{mean},\max\}$. We find that the agents’ learned policies recover the exact heterogeneity gains derived in the theory (Fig. 2). Next, we transfer the same reward structures into embodied, time-extended environments: Multi-goal-capture, 2v2 tag, and football. We show that our curvature theory continues to be and informative in these settings as well. We discuss the learning dynamics that result, and perform further experiments highlighting the difference between neural and behavioral heterogeneity (Bettini et al., 2023), important for understanding our insights. Finally, to study HetGPS, we parametrize the reward structure of Multi-goal-capture using either parametrized Softmax or Power-Sum aggregators (App. I), and run HetGPS to learn parameterizations that maximize the heterogeneity gain. HetGPS learns the theoretically optimal aggregator instantiations, validating its effectiveness at discovering heterogeneous missions. Implementation details and visualizations are available in App. C and App. K.
(i) Task Allocation. We consider a one-step observationless matrix game where $N$ agents need to choose between $M$ tasks. Their actions are effort allocations $r_{ij}$ with $r_{ij}\geq 0$$\sum_{j}r_{ij}=1$, composing matrix $\mathbf{A}$. With aggregators taken from the set $U,T\in\{\min,\mathrm{mean},\max\}$, our goal is to empirically confirm the heterogeneity gains derived in the theory in a learning context. Each time the game is played, all agents obtain the global reward $R(\mathbf{A})$ computed through the double aggregator. We consider two setups: (1) Continuous ($r_{ij}\in\mathbb{R}_{0\leqslant x\leqslant 1}$): agents can distribute their efforts across tasks, (2) Discrete ($r_{ij}\in\{0,1\}$): agents choose only one task. We train with $N=M=4$ for 12 million steps. Fig. 6 shows the evolution of the heterogeneity gains. The final results match exactly the theoretical predictions of Fig. 2 and our curvature theory: concave outer and convex inner aggregators favor heterogeneity. Additional details and results, e.g., for $N=M=2,8$, are in App. J.
Figure 3: Heterogeneity gain for the discrete and continuous matrix games with $N=M=4$ over training iterations. We report mean and standard deviation after 12M frames over 9 random seeds. The final results match the theoretical predictions in the Table of Fig. 2. Solid lines indicate reward structures predicted by theory to have $\Delta R>0$ in either the discrete or continuous setting; dashed lines indicate predicted no gain in both settings.
@@
Figure 4: Heterogeneity gain for Multi-goal-capture and 2v2 Tag throughout training. We report mean and standard deviation for 30 million training frames over 9 random seeds.
@@
(ii-2) 2v2 tag. In our tag experiment, two learning chasers pursue two heuristic escapees in a randomized obstacle field. We define the effort $r_{ij}^{t}$ to be $1$ if chasing agent $i$ manages to capture escaping agent $j$ by time $t$, and 0 otherwise. Whereas Multi-goal-capture had continuous effort allocations, here they are discrete. The global reward is again computed with aggregators $U,T\in\{\min,\text{mean},\max\}$, and is awarded at every time step. This is a sparse reward signal only awarded upon mission success. For example, $U=\min,T=\max$ pays out only if both escapees are each caught by a chaser, encouraging heterogeneity. Training outcomes are summarized in 4(b). We again see that our theoretical results in Fig. 2 (discrete efforts) predict exactly which aggregators will exhibit $\Delta R>0$. More details and visuals are available in App. L and here.
@@ 6 Discussion
Acknowledgements and Disclosure of Funding
@@ Appendix B Code and Data Availability
We attach the code in the supplementary materials. The code contains instructions on how to reproduce the experiments in the paper and dedicated YAML files containing the hyperparameters for each experiment presented. The YAML files are structured according to the HYDRA (Yadan, 2019) framework which allows smooth reproduction as well as systematic and standardized configuration. We further attach all scripts to reproduce the plots in the paper from the experiment results.
@@ Appendix F Sum-Form Aggregators
Lemma F.1 (Schur Properties of Sum-Form Aggregators (Peajcariaac & Tong, 1992)).
@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
J.4 $N=M=8$
To test scalability, we further report results for the discrete matrix game with $N=M=8$ in Tab. 9. As discussed in Sec. 5, due to the larger agent scale and action dimensionality (8 agents and 8 tasks), for some reward structures the empirical heterogeneity gain is negative, since the neurally heterogeneous agents we train require more time to discover the optimal policy, hence lag behind their homogeneous counterparts under a fixed training budget. Increasing the number of training steps steers the negative $\Delta R$ values to 0. This nuance aside, what is important is that the results still follow our theory exactly in terms of which reward structures yield a positive heterogeneity gain. For such reward structures, our theory also precisely predicts the numerical value of $\Delta R$ (Fig. 2) .
@@
Our results, shown in Fig. 4(a), show that our curvature theory reliably predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. The heterogeneity gain is smaller in the latter case because learning dynamics matter: with $U=\min,\,T=\max$ the best homogeneous policy is unique (every agent must steer to the midpoint between the two goals) so homogeneous learners seldom find it, leaving room for heterogeneous policies to excel (see App. K). By contrast, $U=\mathrm{mean},\,T=\max$ admits a continuum of good homogeneous policies, which homogeneous teams execute more easily. For $U=\max,\,T=\min$ and $U=\max,\,T=\mathrm{mean}$, the theoretical $\Delta R$ is 0, yet the empirical heterogeneity gap is negative (4(a)). This occurs because the reward peaks only when all agents coordinate on the same goal. Neurally heterogeneous teams learn this uniform behavior slower than homogeneous teams, so they underperform within the fixed training budget. Additional training would close this gap to $\Delta R=0$.
@@
The goal of our tag experiment is to showcase that our theoretical results, which predict the value of $\Delta R$ based on the curvature of the aggregators, hold for discrete, sparse rewards. Specifically, our results for discrete efforts in Fig. 2 predict that only $(U,T)=(\min,\max),(\min,\text{mean}),(\text{mean},\max)$ will have positive heterogeneity gain, with $(\min,\max)$ maximizing the gain. We show in 4(b) that this holds in 2v2 tag, despite the fact that this is a challenging, embodied, long-horizon, whereas our formal results are for instantaneous allocation games [^4] The gain for $(\text{mean},\max)$ is small compared to the other two aggregator combinations, but still positive at $\Delta R\approx 0.37$. This is also significantly higher than aggregator combinations for which we predict $\Delta R$ is not positive, the largest of which attained $\Delta R<0.01$.. Note that this is a highly interpretable result: a $(\min,\max)$ means that agents are only awarded when both escapers are caught, incentivizing heterogeneous strategies where chaser agents split their behaviour so that the chasing efforts $r^{t}_{ij}$ are equally distributed between both escapers. Fig. 8 visualizes the trajectories learned by agents trained under this $(\min,\max)$ reward structure, showing distinct emergent pursuit strategies emerging depending on whether the agents are neurally heterogeneous or neurally homogeneous.
@@
$U=\text{mean},\;T=\text{mean}$ $0.009\pm 0.075$ No Agents should attend both the opponent and the ball.
@@
We report heterogeneity gains after training homogeneous and heterogeneous policies in Tab. 10. This shows that our curvature test predicts the heterogeneity gain of different reward structures, despite only being a component in the overall reward structure. This insight is important, as it indicates our theoretical insights (the curvature test) may extend beyond environments that strictly follow our task allocation setting.
@@
HetGPS can be formalized as a Stackelberg game, a hierarchical optimization problem involving a leader and followers (Simaan & Cruz, 1973). In our setting:

### v4

# arXiv:2506.09434v4[cs.MA] 01 Mar 2026

@@ Abstract
00footnotetext: Website and code: https://sites.google.com/view/hetgps
@@ 1 Introduction
Theoretical Insights. We first study a pure, non-spatial and instantaneous variant of multi-agent task allocation: each agent commits its effort allocation once, and the team is rewarded immediately (Sec. 2). We start from the observation that team reward in many effort–allocation problems can be expressed as $R(\mathbf{A})=U\bigl(T_{1}(\mathbf{a}_{1}),\dots,T_{M}(\mathbf{a}_{M})\bigr),$ where $\mathbf{A}=(r_{ij})$ is the $N\times M$ matrix of agent effort allocations, and $a_{i}$ is the effort allocation vector of agent $i$. The inner, task-level operator $T_{i}$ assigns a score corresponding to the $N$ agents’ efforts on the $i$th task and the outer operator $U$ combines the resulting $M$ task scores into a scalar global reward. Choosing $T$ and $U$ to be the sum operator $\sum$ recovers the $\sum_{j}\sum_{i}r_{ij}$ reward common in RL, whereas alternatives such as max, min, power means, or soft-max encode very different effort–reward relationships. Assuming such a reward structure, we compare the optimal heterogeneous reward, $R_{\mathrm{het}}$, with the best reward attainable under a homogeneous allocation, $R_{\mathrm{hom}}$, and define their difference as the heterogeneity gain $\Delta R=R_{\mathrm{het}}-R_{\mathrm{hom}}$ (Fig. 1). Our main insight is that $\Delta R$ is determined by the curvature of $T$ and $U$: specifically, whether they are Schur-convex or Schur-concave. These criteria immediately enable us to characterize the heterogeneity gain of broad families of reward functions (Table 3); for instance, the soft-max operator switches from Schur-concave to Schur-convex as its temperature increases. We also find exact expressions for $\Delta R$ in several important cases. These results help explain, for example, why a reward structure that involves a $\min$ operator (usually used to enforce that only one agent should pursue a goal) will require behavioral diversity from the agents (Bettini et al., 2024). We relate our findings to multi-agent reinforcement learning (MARL), where environments may be embodied and time-extended, by setting $R(\mathbf{A_{t}})$ as the stepwise reward over an allocation sequence $(A_{t})_{t=1,\ldots,T}$.
@@
Algorithmic Search. To study heterogeneity in MARL settings not covered by our theoretical analysis, we develop Heterogeneity Gain Parameter Search (HetGPS), a gradient-based algorithm that optimizes parameters $\theta$ of underspecified, differentiable MARL environments via backpropagation to find configurations that maximize or minimize the empirical $\Delta R$ (we assume differentiability for training efficiency, but consider non-differentiable environments in App. P). While HetGPS can in principle optimize any differentiable environment feature to influence $\Delta R$, we use it here to explore reward structures, as a means of verifying and extending our theoretical insights. Maximizing the heterogeneity gain allows us to discover reward functions where behavioral diversity is essential. Minimizing the gain leads us to settings where homogeneous policies are sufficient.
@@
Experiments. We validate our theoretical insights, and HetGPS, in simulation, by evaluating in both single-shot and long-horizon reinforcement learning environments whose reward structure instantiates the kinds of aggregation operators studied. First, in a continuous and a discrete matrix game, we test reward structures based on all nine possible combinations of $\{\min,\mathrm{mean},\max\}$, and find that the heterogeneity gains that result from the agents’ learned policies match our theoretical predictions: concave outer operators and convex inner operators benefit heterogeneous teams. Next, we test the same operators in embodied, partially observable environments: Multi-goal-capture, Tag, and Football. We find that our theory also transfers to such long-horizon MARL settings, and show that reward structures that maximize heterogeneity are meaningful and practically useful. Finally, we find that the empirical heterogeneity gain disappears as the richness of agents’ observations is increased, recovering the finding that rich observations allow agents with identical policy networks to be behaviorally heterogeneous (Bettini et al., 2023; Leibo et al., 2019).
@@
We then turn to HetGPS. Across two parameterizable families of operators (Softmax and Power-Sum), we show that, despite running on embodied environments, HetGPS rediscovers the reward regimes predicted by our curvature theory to maximize the heterogeneity gain, validating both HetGPS and the connection between our theoretical insights and MARL reward design (Sec. 5).
@@ 1.1 Related Works
Behavioral Diversity in MARL. Behavioral heterogeneity, where capability-identical agents learn distinct policies, can markedly improve exploration, robustness, and reward (Bettini et al., 2023). Yet heterogeneity reduces parameter sharing and thus sample-efficiency, so a core practical question is when its benefits outweigh that cost. Existing MARL methods typically adopt one of two poles: endowing each agent with its own network, or enforcing parameter sharing so all agents follow a single policy (Gupta et al., 2017a; Rashid et al., 2020; Foerster et al., 2018; Kortvelesy and Prorok, 2022; Sukhbaatar et al., 2016). A large body of work explores the efficiency–diversity trade-off (Christianos et al., 2021; Fu et al., 2022) by interpolating between these extremes: e.g., injecting agent IDs into the observation (Foerster et al., 2016; Gupta et al., 2017a), masking different subsets of shared weights (Li et al., 2024b), sharing only selected layers (Li et al., 2021), pruning a shared network into agent-specific sub-graphs (Kim and Sung, 2023), or producing per-agent parameters with a hypernetwork (Tessera et al., 2024). Further, several methods for promoting behavioral diversity in MARL have been proposed, such as: conditioning agents’ policies on a latent representation (Wang et al., 2020), decomposing and clustering action spaces (Wang et al., 2021), dynamically grouping agents to share parameters (Yang et al., 2022), applying structural constraints to the agents’ policies (Bettini et al., 2024), or by intrinsic rewards that maximize diversity (Li et al., 2021; Jaques et al., 2019; Wang et al., 2019; Jiang and Lu, 2021; Mahajan et al., 2019; Liu et al., 2023; 2024; Li et al., 2024a). While these studies demonstrate how to obtain diversity, they presume tasks where heterogeneity is advantageous. Our work addresses the orthogonal question of when diversity is beneficial, giving a principled characterization of which reward structures create that incentive in the first place.
@@ 3 Analysis
Focusing on continuous allocations, we ask what properties of aggregators guarantee $\Delta R>0$. We draw on the concept of Schur-convexity. Schur-convex functions can be understood as generalizing symmetric, convex aggregators: every convex and symmetric function is Schur-convex, but a Schur-convex function is not necessarily convex (Roberts and Varberg, 1974; Peajcariaac and Tong, 1992). Proofs for all results are available in App. G.
@@
It is crucial to note that this constant-sum assumption is specific to Thm. 3.3 and is not required for our other results, which apply broadly.
@@ 4 Heterogeneity Gain Parameter Search (HetGPS)
In complex scenarios where theory might be less applicable, we study heterogeneity through algorithmic search. We consider the setting of a Parametrized Dec-POMDP (PDec-POMDP, defined in App. O). A PDec-POMDP represents a Dec-POMDP (Oliehoek et al., 2016), where the observations, transitions, or reward are conditioned on parameters $\theta$. Hence, the return obtained by the agents, $G^{\theta}(\mathbf{\pi})$, can be differentiated with respect to $\theta$: $\nabla_{\theta}G^{\theta}(\mathbf{\pi})=\frac{\partial}{\partial_{\theta}}G^{\theta}(\mathbf{\pi})$. In particular, computing this gradient in a differentiable simulator allows us to back-propagate through time and optimize $\theta$ via gradient ascent [^3] Although the same approach can train policies (Xu et al., 2022; Song et al., 2024), HetGPS instead optimizes environment parameters and policies separately, using standard zeroth-order policy-gradient methods, to avoid being trapped in local minima..
@@ 5 Experiments
To empirically ground our theoretical analysis, we conduct a three-stage experimental study in cooperative MARL. We first analyze a one-step, observation-free matrix game in which each agent allocates effort $r_{ij}$ over $M$ tasks, and consider reward structures defined by aggregator pairs $U,T\in\{\min,\mathrm{mean},\max\}$. We find that the agents’ learned policies recover the exact heterogeneity gains derived in the theory (Fig. 2). Next, we transfer the same reward structures into embodied, time-extended environments: Multi-goal-capture, 2v2 Tag, and Football. We show that our curvature theory continues to be informative in these settings. We discuss the learning dynamics that result, and perform further experiments highlighting the difference between neural and behavioral heterogeneity (Bettini et al., 2023), important for understanding our insights. Finally, to study HetGPS, we parametrize the reward structure of Multi-goal-capture using either parametrized Softmax or Power-Sum aggregators (App. I), and run HetGPS to learn parameterizations that maximize the heterogeneity gain. HetGPS learns the theoretically optimal aggregator instantiations, validating its effectiveness at discovering heterogeneous missions. Implementation details and visualizations are available in App. C and App. K.
(i) Task Allocation. We consider a one-step observationless matrix game where $N$ agents need to choose between $M$ tasks. Their actions are effort allocations $r_{ij}$ with $r_{ij}\geq 0$$\sum_{j}r_{ij}=1$, composing matrix $\mathbf{A}$. With aggregators taken from the set $U,T\in\{\min,\mathrm{mean},\max\}$, our goal is to empirically confirm the heterogeneity gains derived in the theory in a learning context. Each time the game is played, all agents obtain the global reward $R(\mathbf{A})$ computed through the double aggregator. We consider two setups: (1) Continuous ($r_{ij}\in\mathbb{R}_{0\leqslant x\leqslant 1}$): agents can distribute their efforts across tasks, (2) Discrete ($r_{ij}\in\{0,1\}$): agents choose only one task. We train with $N=M=4$ for 12 million steps. Fig. 6 shows the evolution of the heterogeneity gains. The final results match exactly the theoretical predictions of Fig. 2 and our curvature theory: concave outer and convex inner aggregators favor heterogeneity. Additional details and results, e.g., for $N,M\in\{2,8,11\}$, are in App. J.
Figure 3: Heterogeneity gain for the discrete and continuous matrix games with $N=M=4$ over training iterations. We report mean and standard deviation after 12M frames over 9 random seeds. The final results match the theoretical predictions in the Table of Fig. 2. Solid lines indicate reward structures predicted by theory to have $\Delta R>0$ in either the discrete or continuous setting; dashed lines indicate predicted no gain in both settings. Final gain values are reported in Tab. 7 and Tab. 8.
@@
Figure 4: Heterogeneity gain for Multi-goal-capture and 2v2 Tag throughout training. We report mean and standard deviation for 30 million training frames over 9 random seeds. Final gain values are reported in Tab. 11 and Tab. 12.
@@
(ii-2) 2v2 Tag. In our tag experiment, two learning chasers pursue two heuristic escapees in a randomized obstacle field. We define the effort $r_{ij}^{t}$ to be $1$ if chasing agent $i$ manages to capture escaping agent $j$ by time $t$, and 0 otherwise. Whereas Multi-goal-capture had continuous effort allocations, here they are discrete. The global reward is again computed with aggregators $U,T\in\{\min,\text{mean},\max\}$, and is awarded at every time step. This is a sparse reward signal only awarded upon mission success. For example, $U=\min,T=\max$ pays out only if both escapees are each caught by a chaser, encouraging heterogeneity. Training outcomes are summarized in 4(b). We again see that our theoretical results in Fig. 2 (discrete efforts) predict exactly which aggregators will exhibit $\Delta R>0$. More details, visuals, and experiments with greater number of agents are available in App. L and here.
@@ 6 Discussion
Reproducibility Statement
Our supplementary website contains the source code we used to produce all results in this work, including the code used to train the agents, our implementation of HetGPS, and the code used to produce all plots in the paper (see Appendix C). The readme contains detailed instructions on how to use this code. All mathematical claims made in this work are fully proven in the Appendix, and our assumptions are described in detail in Section 2 (“Problem Setting”). Appendices Q and P address potential questions readers may have regarding the stability of the bilevel optimization process used in HetGPS and similar environment design algorithms in the literature. Finally, Appendix R addresses the assumptions and scope of our work, and outlines some remaining open questions.
Acknowledgments
@@ Appendix B Code and Data Availability
We provide the link to the code in the supplementary website. The code contains instructions on how to reproduce the experiments in the paper and dedicated YAML files containing the hyperparameters for each experiment presented. The YAML files are structured according to the HYDRA (Yadan, 2019) framework which allows smooth reproduction as well as systematic and standardized configuration. We further attach all scripts to reproduce the plots in the paper from the experiment results.
@@ Appendix F Sum-Form Aggregators
Lemma F.1 (Schur Properties of Sum-Form Aggregators (Peajcariaac and Tong, 1992)).
@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
J.4 $N=M=8$ and $N=12,M=2$
To test scalability, we further report results for the discrete matrix game with $N=M=8$ in Tab. 9, $N=11,M=2$ in Tab. 10. In the case of $N=11,M=2$, results match the predictions of Table 2 (discrete rewards) precisely.
For $N=M=8$, we match the predictions precisely except for two cells which have negative empirical $\Delta R$. As discussed in Sec. 5, due to the larger agent scale and action dimensionality (8 agents and 8 tasks), for some reward structures the empirical heterogeneity gain is negative, since the neurally heterogeneous agents we train require more time to discover the optimal policy, hence lag behind their homogeneous counterparts under a fixed training budget. Increasing the number of training steps steers the negative $\Delta R$ values to 0. This nuance aside, what is important is that the results still follow our theory exactly in terms of which reward structures yield a positive heterogeneity gain. For such reward structures, our theory also precisely predicts the numerical value of $\Delta R$ (Fig. 2) .
@@
Max $75\pm 0.5 0.0$$\pm 0.0 0.0$$0\pm 0.0$
Table 10: Heterogeneity gain $\Delta R\in\mathbb{R}_{0\leqslant x\leqslant 1}$ of the discrete matrix game with $N=11,M=2$. We report mean and standard deviation after 12 million training frames over 8 different random seeds. The results match the predictions of Table 2 (discrete rewards) precisely.
$T$
Min Mean Max
$U$ Min $0.0\pm 0.0$ $\bm{0.454545\approx 5/11}\pm 0.0$ $\bm{1.0}\pm 0.0$
Mean $94\pm 0.0 0.0±0.0$$\pm 0.0$ $\bm{0.5}\pm 0.0$
@@
Our results, shown in 4(a) and Tab. 11, show that our curvature theory reliably predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. The heterogeneity gain is smaller in the latter case because learning dynamics matter: with $U=\min,\,T=\max$ the best homogeneous policy is unique (every agent must steer to the midpoint between the two goals) so homogeneous learners seldom find it, leaving room for heterogeneous policies to excel (see App. K). By contrast, $U=\mathrm{mean},\,T=\max$ admits a continuum of good homogeneous policies, which homogeneous teams execute more easily. For $U=\max,\,T=\min$ and $U=\max,\,T=\mathrm{mean}$, the theoretical $\Delta R$ is 0, yet the empirical heterogeneity gap is negative (4(a)). This occurs because the reward peaks only when all agents coordinate on the same goal. Neurally heterogeneous teams learn this uniform behavior slower than homogeneous teams, so they underperform within the fixed training budget. Additional training would close this gap to $\Delta R=0$.
Table 11: Heterogeneity gain at the end of training for the Multi-goal-capture experiments in 4(a). We report mean and standard deviation after 30 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $0\pm 0.02 0.0$$01\pm 0.02$ $\bm{0.21}\pm 0.03$
Mean $3\pm 0.01 0.0±0.0$$\pm 0.0$ $\bm{0.12}\pm 0.09$
Max $08\pm 0.06 0.0$$7\pm 0.08 0.0$$0\pm 0.0$
@@
The goal of our tag experiment is to showcase that our theoretical results, which predict the value of $\Delta R$ based on the curvature of the aggregators, hold for discrete, sparse rewards. Specifically, our results for discrete efforts in Fig. 2 predict that only $(U,T)=(\min,\max),(\min,\text{mean}),(\text{mean},\max)$ will have positive heterogeneity gain, with $(\min,\max)$ maximizing the gain. We show in 4(b) and Tab. 12 that this holds in 2v2 tag, despite the fact that this is a challenging, embodied, long-horizon, whereas our formal results are for instantaneous allocation games [^4] The gain for $(\text{mean},\max)$ is small compared to the other two aggregator combinations, but still positive at $\Delta R\approx 0.37$. This is also significantly higher than aggregator combinations for which we predict $\Delta R$ is not positive, the largest of which attained $\Delta R<0.01$.. Note that this is a highly interpretable result: a $(\min,\max)$ means that agents are only awarded when both escapers are caught, incentivizing heterogeneous strategies where chaser agents split their behaviour so that the chasing efforts $r^{t}_{ij}$ are equally distributed between both escapers. Fig. 8 visualizes the trajectories learned by agents trained under this $(\min,\max)$ reward structure, showing distinct emergent pursuit strategies emerging depending on whether the agents are neurally heterogeneous or neurally homogeneous.
To test the robustness of our predictions to greater number of agents, we also ran an experiment with $11$ agents: $8$ chasers and $3$ escapers. We trained the agents over $500$ episodes of length $1000$ each (this episode length is more than twice as long as our other experiments, indicating that our predictions are stable over longer horizons). Note that we still collect the same amount of total frames (30M) as we reduce the number of environments sampled in parallel.
Due to the high computational cost associated with these experiments, we selected $2$ aggregator combinations for which Table 2 predicts a positive $\Delta R$: $(U,T)=(\textrm{min},\textrm{mean})$ and $(U,T)=(\textrm{min},\textrm{max})$, and we further selected three “control” combinations for which we expect $\Delta R\leq 0$. The results, shown in Figure 9 (discrete rewards), illustrate that our predictions still hold in this case. The final $\Delta R$ values are $(U=\text{min},T=\text{mean})=1.243\pm 0.615$, $(U=\text{min},T=\text{max})=0.112\pm 0.084$, $(U=\text{mean},T=\text{mean})=-0.196\pm 0.053$, $(U=\text{mean},T=\text{max})=-0.990\pm 1.025$, and $(U=\text{max},T=\text{max})=-3.709\pm 2.491$.
It is important to note that, while our theoretical predictions regarding when $\Delta R>0$ hold for any number of agents $N$, they specifically tell us what happens when agents allocate their efforts optimally. The empirical heterogeneity gain $\Delta R$ crucially depends on the quality of strategy agents learn in practice, which in turn also depends on the transition dynamics of the environment, complicating things. When the number of agents or complexity of the task is very large, we may eventually witness a divergence between the empirical heterogeneity gain and the theoretical predictions, for this reason. This does not indicate a problem with our theoretical predictions. Rather, it is a limitation of learning-based methods; using better methods in environments which enable optimal effort allocation will lead to empirical results that more closely mirror our predictions.
Figure 9: Heterogeneity gain for 8v3 Tag throughout training. We report mean and standard deviation for 30 million training frames over 8 random seeds. Positive aggregator combinations are colored and follow the predictions of Table 2 for discrete rewards. Final gain values are reported in the text of Appendix L.
Table 12: Heterogeneity gain at the end of training for the Tag experiments in 4(b). We report mean and standard deviation after 30 million training frames over 9 different random seeds.
$T$
Min Mean Max
$U$ Min $0.0\pm 0.0$ $\bm{1.68}\pm 0.24$ $\bm{3.47}\pm 0.23$
Mean $03\pm 0.04 0.0$$02\pm 0.06$ $\bm{0.36}\pm 1.44$
Max $0.02\pm 0.09$ $0.11\pm 0.18$ $0.30\pm 0.15$
@@
$U=\text{mean},\;T=\text{mean}$ $0.01\pm 0.07$ No Agents should attend both the opponent and the ball.
@@
We report heterogeneity gains after training homogeneous and heterogeneous policies in Tab. 13. This shows that our curvature test predicts the heterogeneity gain of different reward structures, despite only being a component in the overall reward structure. This insight is important, as it may indicative our theoretical insights (the curvature test) may extend beyond environments that strictly follow our task allocation setting. However, this is just one possible scenario, and as such, this possibility requires further verification in future work.
@@
HetGPS can be formalized as a Stackelberg game, a hierarchical optimization problem involving a leader and followers (Simaan and Cruz, 1973). In our setting: